• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name D. Lambert

    Log entry time 10:27:03 on June25,2004

    Entry number 126817

    keyword=IOC reboot (again)

    had to reboot IOC at about 10:20.  We had regained beam a minute or two before, but seemed unrelated to beam.  ESR flow flat-lined first, white alarm occurred late (after I'd started reboot process).  Lost beam during heater-off period, so had to really push the manual controls, but all is good now.  strange that it's happened twice in two hours.